OpenStudy (anonymous):

can you solve (x+y)dx+(x-y)dy=0 please , exam in few hours

OpenStudy (amistre64):

cant recall the details; but something about integrating one side and then deriving it down the other .....

OpenStudy (amistre64):

M(x,y) + N(x,y) .... and some such

OpenStudy (amistre64):

int(x+y)dx = (x^2)/2 + xy + g(y)

OpenStudy (anonymous):

Well this is how I would approach it: \[(x-y)dy=-(x+y)dx \rightarrow (x-y)\frac{dy}{dx}=-(x+y)\] From here \[\frac{dy}{dx}=\frac{-(x+y)}{x-y}\] Does that help?

OpenStudy (amistre64):

Dy((x^2)/2 + xy + g(y)) = x + g'(y)

OpenStudy (anonymous):

thank you very much

OpenStudy (amistre64):

g'(y) = -y; so g(y) = -(y^2)/2 + C maybe? to get us to: (x^2)/2 + xy -(y^2)/2 + C ?

OpenStudy (anonymous):

ok thx i'll try to approach it this way

OpenStudy (anonymous):

thx you all

OpenStudy (amistre64):

\(\color{red}{\text{youre welcome}}\) \(\color{magenta}{\text{:)}}\)